21/2/22 16:00
Call from main control room, beam is ON, probably has been for several hours judging by rate on xy total plot.
Changes made to online monitoring code by OH, include setting to only 1 DSSD, correct offsets from the pulser walkthrough, offline analysis capabilities and X/Y position correction.
Monitoring code restarted in CARME_SORT/build_no_output
17:31 Run 50 ended, run 51 started now beam is definitely in the ring. Tried to insert span zero schottky into aida2 correlation scalar, however rate was far too high.
18:03 Removed span zero schottky signal, started new run 52.
With beam ON we noticed some missing channels in the xy total plot. Aida 1 and 2 stat totals on spectrum browser show missing channels. Asic check on both to no effect. Fees power cycled.
19:57 Started run 53. Currently waiting for beam
21:20 Beam ON
1*L plots checked for aida 1, several channels showed no data (noise)
Slow comparator LEC threshold lowered to 0x32 to aida 1, checked 1*L plots and channels are now showing lots of noise. So those channels appear to be quiet using 0x64 threshold. (attachment 3)
Threshold increased back to 0x64
22:33
System checks OK -white rabbit fails on all. ~30 seperation
Temperatures OK
Statistics within range - new scalars added for aida02 and aida05
Data being written ~300 kb/sec
Leakage current ~3.6 uA
00:39
System checks checks OK -white rabbit fails on all. ~42 seperation
Temperatures OK
Statistics within range with new scalars
Data being written 256 kb/sec
Leakage current ~3.615 uA
XY counts dominated by hot channels - most blue pixels are ~10 counts
|